# mGTE: Generalized Long-Context Text Representation and Reranking Models for Multilingual Text Retrieval

### v1

@@ C.2 GLUE
We evaluate the GLUE benchmark based on the scripts272727 github.com/huggingface/transformers/tree/main/examples/pytorch/text-classification#glue-tasks and data282828 hf.co/datasets/nyu-mll/glue provided by transformers. In fine-tuning of each dataset, we use the epoch number of 3, learning rate of 2e-5, batch size of 32, and max sequence length of 128. For MRPC, STS-B, and RTE, we start from the checkpoint of MNLI following Liu et al. (2019). The MNLI checkpoint is shared with XNLI of XTREME-R (§C.1).
@@ Appendix D Text Embedding Evaluation
Table 14: Results on MTEB English subset Muennighoff et al. (2023). We compare models from the online leaderboard, where derived models (developed by secondary training on other public off-the-shelf models) are not listed. Denote our runs. To be consistent with the setting in contrastive pre-training, in retrieval tasks, the max sequence length of the document side is set to 1024.
@@
Table 14 shows the results on English MTEB Muennighoff et al. (2023). For reference, we include our English embedding models (Our-en-base/large-embed, trained by the two-stage contrastive learning on the English part of our data) and top-performing systems from the online leaderboard. We can see that the multilingual models still have a noticeable gap compared to the English models.

### v2

@@ C.2 GLUE
We evaluate the GLUE benchmark based on the scripts292929 github.com/huggingface/transformers/tree/main/examples/pytorch/text-classification#glue-tasks and data303030 hf.co/datasets/nyu-mll/glue provided by transformers. In fine-tuning of each dataset, we use the epoch number of 3, learning rate of 2e-5, batch size of 32, and max sequence length of 128. For MRPC, STS-B, and RTE, we start from the checkpoint of MNLI following Liu et al. (2019). The MNLI checkpoint is shared with XNLI of XTREME-R (§C.1).
@@ Appendix D Text Embedding Evaluation
Table 14: Results on MTEB English subset Muennighoff et al. (2023a). We compare models from the online leaderboard, where derived models (developed by secondary training on other public off-the-shelf models) are not listed. Denote our runs. To be consistent with the setting in contrastive pre-training, in retrieval tasks, the max sequence length of the document side is set to 1024.
@@
Table 14 shows the results on English MTEB Muennighoff et al. (2023a). For reference, we include our English embedding models (Our-en-base/large-embed, trained by the two-stage contrastive learning on the English part of our data) and top-performing systems from the online leaderboard. We can see that the multilingual models still have a noticeable gap compared to the English models.